Quantitative Problem: Adams Manufacturing Inc. buys $10.4 million of materials (net of discounts) on terms of 2/10, net 60; and it currently pays after 10 days and takes the discounts. Adams plans to expand, which will require additional financing. If Adams decides to forgo discounts, how much additional credit could it obtain? Round your answer to the nearest cent. Do not round your intermediate calculations. Use 365 day in a year.
$

What would be the nominal and effective cost of such a credit? Round your answer to 2 decimal places. Do not round intermediate calculations. Use 365 day in a year.
Nominal cost:   %
Effective cost:   %

If the company could receive the funds from a bank at a rate of 7.3%, interest paid monthly, based on a 365-day year, what would be the effective cost of the bank loan? Round your answer to 2 decimal places. Do not round intermediate calculations.
%

please Help!Thank you!
Answer #1

Answer:-

a) Purchases = $10,400,000

terms = 2/10 net 60

currently pays after Day 10 and takes discounts.

If Adams forgoes the discount, then additional credit is (10,400,000/365)×(60-10)

=1,424,657.53 and has to pay within 60 days.

b) Nominal cost of credit = [discount percent/ (1 - discount percent)] × [365 / (full payment days - discount days)]

Nominal cost = [2%/ (1 - 2%)] × [365 / (60 - 10)]

Nominal cost = 14.89%

Effective cost of trade credit = [1 + (2/100-2)] 365/(60-10) - 1

Effective Cost = (1 + 0.0204)7.3 -1

Effective Cost = 1.1588 - 1 = 0.1588 or 15.88%

c) Bank rate = 7.3%

Effective annual rate =(1 + 0.073/12) 12 −1 =7.23%

Effective annual rate =1.07549 - 1

Effective annual rate = 0.07549 or 7.549%

​​